# Benford behavior resulting from stick and box fragmentation processes

### v2

@@ 2.1. Preliminary
In [FILM], Fang, Irons, Lippelman and Miller proved that if $y_{i}\in\mathbb{Q}$ for all $1\leq i\leq m-1$, then the distribution of the stick lengths does not follow Benford behavior as $N\rightarrow\infty$. To be exact, they were able to show that the distribution of the logarithms of the stick lengths converges to a discrete distribution, contrary to a uniform distribution on $[0,1]$ which is an equivalent condition for convergence to strong Benford behavior by Uniform Distribution Characterization 1.4. Hence, having $y_{i}\not\in\mathbb{Q}$ for some $1\leq i\leq m-1$ is a necessary condition for the distribution of the stick lengths to follow Benford behavior.
@@
For the remainder of this section, we show that having $y_{i}\not\in\mathbb{Q}$ for some $1\leq i\leq m-1$ is also a sufficient condition for the distribution of the stick lengths to converge to Benford’s behavior. Since we can always reorder the $p_{i}^{k_{i}}$’s in the stick length $A_{k_{1},k_{2},\dots,k_{m}}:=p_{1}^{k_{1}}p_{2}^{k_{2}}\cdots p_{m}^{k_{m}}$ and thus the order of $y_{i}$’s in the factorization, then it suffices to show that the first exponent $y_{1}\not\in\mathbb{Q}$ is a sufficient condition for the distribution of the stick lengths to converge to Benford’s behavior.
We know that after stage $N$, the number of sticks of length $A_{k_{1},k_{2},\dots,k_{m}}$ is equal to the multinomial coefficient
$\binom{k_{1}+k_{2}+\cdots+k_{m}}{k_{1},k_{2},\dots,k_{m}}\ =\ \binom{N}{k_{1},k_{2},\dots,k_{m}}$. (2.1)
@@
From the proof of Theorem , we know that the stick length $A_{k_{1},k_{2},\dots,k_{m}}$ has the factorization
$A_{k_{1},k_{2},\dots,k_{m}}\ =\ p_{1}^{k_{1}}p_{2}^{k_{2}}\cdots p_{m}^{k_{m}}\ =\ \left(\frac{p_{1}}{p_{2}}\right)^{k_{1}}\left(\frac{p_{2}}{p_{3}}\right)^{k_{1}+k_{2}}\cdots\left(\frac{p_{m-1}}{p_{m}}\right)^{\sum_{j=1}^{m-1}k_{j}}(p_{m})^{N},$ (2.2)
which motivates us to define the exponents $y_{i}=\log(p_{i}/p_{i+1})$ for $1\leq i\leq m-1$. Now, fix an interval $(a,b)\subset(0,1)$. Let
$\chi_{(a,b)}(x)\ :=\ \mathbbm{1}\Big{(}\log(x)\textup{ mod 1}\in(a,b)\Big{)}$ (2.3)
@@
be the indicator function for $\log(x)\textup{ mod 1}\in(a,b)$. Then after stage $N$, the probability that a stick length mod 1 is in $(a,b)$ equals
@@
$F_{N}(a,b)\ :=\ \frac{1}{m^{N}}\sum_{\begin{subarray}{c}0\leq k_{1},k_{2}\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}\chi_{(a,b)}\left(\prod_{i=1}^{m}p_{i}^{k_{i}}\right).$ (2.4)
@@
$F(a,b)\ :=\ \lim_{N\rightarrow\infty}F_{N}(a,b)\ =\ b-a,$ (2.5)
which is a sufficient condition for Benford behavior.
2.2. Combinatorial identities
The key steps to our proof are two combinatorial identities related to multinomial coefficients, which we present and prove below. We first review the multinomial coefficients.
Definition 2.1.
@@ 2.2. Combinatorial identities
$\binom{n}{k_{1},k_{2},\dots,k_{m}}\ :=\ \frac{n!}{k_{1}!k_{2}!\cdots k_{m}!}.$ (2.6)
The following identity on the factorization of multinomial coefficients allows us to express multinomial coefficients in terms of product of binomial coefficients. As we shall see, this forms the key ingredient that reduces our multi-proportion problem into the original single-proportion problem in [B–].
Lemma 2.2.
[Stan] For $m\geq 1$,
$\binom{k_{1}+k_{2}+\cdots+k_{m}}{k_{1},k_{2},\dots,k_{m}}\ =\ \prod_{i=1}^{m}\binom{k_{1}+\cdots+k_{i}}{k_{i}}$. (2.7)
The proof for the identity above is a straightforward induction through the expansion of the multinomial and binomial coefficients. Below is another identity that is especially useful later for evaluating sums of multinomial coefficients.
Lemma 2.3.
[Stan] For $N\in\mathbb{Z}_{\geq 0}$ and $m\in\mathbb{Z}_{>0}$,
$m^{N}\ =\ \sum_{\begin{subarray}{c}0\leq k_{1},k_{2},\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}\ =\ \sum_{\begin{subarray}{c}0\leq k_{1},\dots,k_{m-1}\leq N\\ k_{1}+k_{2}+\cdots+k_{m-1}=N\end{subarray}}2^{k_{1}}\binom{N}{k_{1},k_{2},\cdots,k_{m-1}}$. (2.8)
The most straightforward proof of the identity above relies on two different ways of representing and expanding $m^{N}$: one as $(\underbrace{1+\cdots+1}_{\textrm{$N$}})^{N}$ and the other as $(2+\underbrace{1+\cdots+1}_{\textrm{$N-2$}})^{N}$. See [Stan] for more.
2.3. Truncation
We prove that for any $\epsilon\in(0,1)$, the contribution to $F_{N}(a,b)$ of multinomial coefficients with $k_{1}+k_{2}<N^{\epsilon}$ is negligible. This allows us to only consider $k_{1}+k_{2}$ that are sufficiently large for these stick lengths $A_{k_{1},k_{2},\dots,k_{m}}$.
Proposition 2.4.
For any $\epsilon\in(0,1)$,
$\sum_{\begin{subarray}{c}0\leq k_{1},k_{2}\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}\ \leq\ (m-2)^{N}N^{N^{\epsilon}+2}.$ (2.9)
@@ 2.3. Truncation
$\ \leq\ N^{k_{1}+k_{2}}\cdot\binom{N-(k_{1}+k_{2})}{k_{3},k_{4},\dots,k_{m}}$. (2.10)
@@
$\sum_{\begin{subarray}{c}0\leq k_{1},k_{2},\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}$ $\ =\ \sum_{\begin{subarray}{c}0\leq k_{1},k_{2}<N^{\epsilon}\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}N^{k_{1}+k_{2}}\sum_{\begin{subarray}{c}0\leq k_{3},k_{4},\dots,k_{m}\leq N-(k_{1}+k_{2})\\ k_{3}+k_{4}+\cdots+k_{m}=N-(k_{1}+k_{2})\end{subarray}}\binom{N-(k_{1}+k_{2})}{k_{3},k_{4},\dots,k_{m}}$
@@
$\ \leq\ (m-2)^{N}(N^{\epsilon})^{2}N^{N^{\epsilon}}\ \leq\ (m-2)^{N}N^{N^{\epsilon}+2}.$ (2.11)
@@
$\frac{1}{m^{N}}\sum_{\begin{subarray}{c}0\leq k_{1},k_{2}\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}$. (2.12)
According to Proposition 2.4, we have that (2.12) is bounded above by $((m-2)/m)^{N}N^{N^{\epsilon}+2}$. Hence, the logarithm of the probability is bounded above by
$\log\left(\left(\frac{m-2}{m}\right)^{N}N^{N^{\epsilon}+2}\right)\ =\ N\log\left(\frac{m-2}{m}\right)+(N^{\epsilon}+2)\log(N).$ (2.13)
It is clear that (2.13) goes to -$\infty$ as $N\rightarrow\infty$. Thus the probability goes to 0 as $N\rightarrow\infty$. So it suffices to consider the case where $k_{1}+k_{2}\geq N^{\epsilon}$. Now, fix $k(N)\geq N^{\epsilon}$ and let $0\leq k_{1},k_{2}\leq k(N)$ with $k_{1}+k_{2}=k(N)$ and $0\leq k_{3},k_{4},\dots,k_{m}\leq N-k(N)$ with $k_{3}+k_{4}+\dots+k_{m}=N-k(N)$. By [B–], we know that the frequency $k_{1}$ of stick length $A_{k_{1},k_{2},\dots,k_{m}}=p_{1}^{k_{1}}p_{2}^{k_{2}}\dots p_{m}^{k_{m}}$ follows a binomial distribution with mean $k(N)/2$ and standard deviation $\sqrt{k(N)}/2$. Pick some $\delta\in(0,\epsilon/10)$. We see that it suffices to consider cases where $|k_{1}-k(N)/2|<(\lceil N^{\delta}\rceil\sqrt{k(N)})/2$, because the probability that $k_{1}$ is outside this range is asymptotically small by Chebyshev’s inequality:
$\mathbb{P}\left(\left|k_{1}-\frac{k(N)}{2}\right|\ \geq\ \frac{\lceil N^{\delta}\rceil\sqrt{k(N)}}{2}\right)\ \leq\ \frac{1}{\lceil N^{\delta}\rceil^{2}}.$ (2.14)
2.4. Near uniform probability within small intervals
@@
$k_{1,\ell}\ :=\ f_{\ell}\left(\frac{k(N)}{2}\right)+\ell\lceil N^{\delta}\rceil,\quad k_{1,\ell,i}\ :=\ f_{\ell}\left(\frac{k(N)}{2}\right)+\ell\lceil N^{\delta}\rceil+i,\quad 0\leq i\leq N^{\delta}-1,$ (2.15)
@@
We follow Section (5.52) of [B–] to obtain a bound for the difference.
Proposition 2.5.
@@
$\left|\binom{k(N)}{k_{1,\ell}}-\binom{k(N)}{k_{1,\ell+1}}\right|\ \leq\ O\left(\binom{k(N)}{k_{1,\ell}}\cdot N^{-\frac{3\epsilon}{10}}\right).$ (2.17)
Since the proof follows similar argument to Section (5.52) of [B–], we leave the proof details to Appendix A.
2.5. Equidistribution within small intervals
For fixed $k(N):=k_{1}+k_{2},k_{3},k_{4}\dots,k_{m}$, and $\ell$, the stick length
$A_{k_{1,\ell,i},k_{2},\dots,k_{m}}\ =\ \left(\frac{p_{1}}{p_{2}}\right)^{k_{1,\ell,i}}\left(\frac{p_{2}}{p_{3}}\right)^{k(N)}\left(\frac{p_{3}}{p_{4}}\right)^{k(N)+k_{3}}\cdots\left(\frac{p_{m-1}}{p_{m}}\right)^{k(N)+\sum_{j=3}^{m-1}k_{j}}(p_{m})^{N}$ (2.18)
@@ 2.5. Equidistribution within small intervals
Definition 2.6.
[KN] Suppose that $x$ is a real number. The irrationality exponent $\mu$ of $x$ is the supremum of the set of $\mu$ such that $0<|x-p/q|<1/q^{\mu}$ is satisfied by an infinite number of coprime integer pairs $(p,q)$ with $q>0$.
Theorem 2.7.
[KN, Theorem 3.2, Example 3.2] Let $\kappa$ be the irrationality exponent of $\log_{B}(p_{1}/p_{2})$. If $\kappa<\infty$, then
$|J_{\ell}(a,b)|\ =\ (b-a)N^{\delta}+O\left(N^{\delta\left(1-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right),$ (2.19)
for some $\epsilon^{\prime}>0$ such that $\frac{1}{\kappa}+\epsilon^{\prime}<0$ and there is a power saving with the error term. If $\kappa=\infty$, then
$|J_{\ell}(a,b)|\ =\ (b-a)N^{\delta}+o(N^{\delta}).$ (2.20)
In either case, the error term is asymptotically much smaller than the main term $(b-a)N^{\delta}$. Thus, if $\log_{B}(p_{1}/p_{2})\not\in\mathbb{Q}$, then $\log_{B}(A_{k_{1,\ell,i},k_{2},\dots,k_{m}})$ converges to being equidistributed mod 1.
2.6. Evaluation of sums within and across intervals

### v3

@@ 2.1 Preliminary
In [13], Fang, Irons, Lippelman and Miller proved that if $y_{i}\in\mathbb{Q}$ for all $1\leq i\leq m-1$, the distribution of the logarithms mod 1 of the stick lengths converges to a discrete distribution instead of a uniform distribution on $[0,1]$. Hence, having $y_{i}\not\in\mathbb{Q}$ for some $1\leq i\leq m-1$ is a necessary condition for the distribution of the stick lengths to follow strong Benford behavior.
@@
For the remainder of this section, we show that having $y_{i}\not\in\mathbb{Q}$ for some $1\leq i\leq m-1$ is also a sufficient condition for the distribution of the stick lengths to converge to strong Benford’s behavior. After reordering the $p_{i}^{k_{i}}$’s and thus the $y_{i}$’s in the factorization of the stick length $A^{(N)}_{k_{1},k_{2},\dots,k_{m}}:=p_{1}^{k_{1}}p_{2}^{k_{2}}\cdots p_{m}^{k_{m}}$, it suffices to show that having the first exponent $y_{1}\not\in\mathbb{Q}$ is a sufficient condition for the distribution of the stick lengths to converge to strong Benford’s behavior.
We know that after stage $N$, there are
$\binom{k_{1}+k_{2}+\cdots+k_{m}}{k_{1},k_{2},\dots,k_{m}}\ =\ \binom{N}{k_{1},k_{2},\dots,k_{m}}$. (11)
@@
number of sticks of length $A^{(N)}_{k_{1},k_{2},\dots,k_{m}}$. From the proof of Theorem 1 of [13], we know that the stick length $A^{(N)}_{k_{1},k_{2},\dots,k_{m}}$ has the factorization
$A^{(N)}_{k_{1},k_{2},\dots,k_{m}}\ =\ p_{1}^{k_{1}}p_{2}^{k_{2}}\cdots p_{m}^{k_{m}}\ =\ \left(\frac{p_{1}}{p_{2}}\right)^{k_{1}}\left(\frac{p_{2}}{p_{3}}\right)^{k_{1}+k_{2}}\cdots\left(\frac{p_{m-1}}{p_{m}}\right)^{\sum_{j=1}^{m-1}k_{j}}(p_{m})^{N},$ (12)
which motivates us to define the exponents $y_{i}=\log_{B}(p_{i}/p_{i+1})$ for $1\leq i\leq m-1$. Now, fix an interval $(a,b)\subset(0,1)$. Let
$\chi_{(a,b)}(x)\ :=\ \mathbbm{1}\Big(\log_{B}(x)\textup{ mod 1}\in(a,b)\Big)$ (13)
@@
be the indicator function for $\log_{B}(x)\textup{ mod 1}\in(a,b)$. Then after stage $N$, the probability that the logarithm mod 1 of a stick length is in $(a,b)$ equals
@@
$F_{N}(a,b)\ :=\ \frac{1}{m^{N}}\sum_{\begin{subarray}{c}0\leq k_{1},k_{2}\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}\chi_{(a,b)}\left(\prod_{i=1}^{m}p_{i}^{k_{i}}\right).$ (14)
@@
$F(a,b)\ :=\ \lim_{N\rightarrow\infty}F_{N}(a,b)\ =\ b-a,$ (15)
which by the Uniform Distribution Characterization implies strong Benford behavior.
2.2 Combinatorial identities
The key steps to our proof are two combinatorial identities related to multinomial coefficients, which we present in this subsection. We first review the definition of multinomial coefficients.
Definition 9.
@@ 2.2 Combinatorial identities
$\binom{n}{k_{1},k_{2},\dots,k_{m}}\ :=\ \frac{n!}{k_{1}!k_{2}!\cdots k_{m}!}.$ (16)
The following identity allows us to express multinomial coefficients as the product of binomial coefficients. As we shall see, this forms a key ingredient to our proof, as it helps reduce our multi-proportion problem to the original single-proportion problem in [1].
Lemma 1.
[20] For $m\geq 1$,
$\binom{k_{1}+k_{2}+\cdots+k_{m}}{k_{1},k_{2},\dots,k_{m}}\ =\ \prod_{i=1}^{m}\binom{k_{1}+\cdots+k_{i}}{k_{i}}$. (17)
The proof for Lemma 1 follows a straightforward inductive argument based on the expansion of the multinomial and binomial coefficients. Below is another identity that comes in handy when evaluating sums of multinomial coefficients.
Lemma 2.
[20] For $N\in\mathbb{Z}_{\geq 0}$ and $m\in\mathbb{Z}_{>0}$,
$m^{N}$ $\ =\ \sum_{\begin{subarray}{c}0\leq k_{1},k_{2},\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}\ =\ \sum_{\begin{subarray}{c}0\leq k_{1},\dots,k_{m-1}\leq N\\ k_{1}+k_{2}+\cdots+k_{m-1}=N\end{subarray}}2^{k_{1}}\binom{N}{k_{1},k_{2},\cdots,k_{m-1}}$. (18)
The proof for Lemma 2 relies on two different representation and expansion of $m^{N}$: one as $(\underbrace{1+\cdots+1}_{\textrm{$N$}})^{N}$ and the other as $(2+\underbrace{1+\cdots+1}_{\textrm{$N-2$}})^{N}$. See [20] for more.
2.3 Truncation
First, we would like to make some truncation to the sum in (14). We show that for any $\epsilon\in(0,1)$, the contribution to $F_{N}(a,b)$ of multinomial coefficients with $k_{1}+k_{2}<N^{\epsilon}$ is negligible. This allows us to only consider stick lengths $A_{k_{1},k_{2},\dots,k_{m}}$ where $k_{1}+k_{2}$ are sufficiently large.
Proposition 2.
For any $0<\epsilon<1$,
$\sum_{\begin{subarray}{c}0\leq k_{1},k_{2}\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}\ \leq\ (m-2)^{N}N^{N^{\epsilon}+2}.$ (19)
@@ 2.3 Truncation
$\ \leq\ N^{k_{1}+k_{2}}\cdot\binom{N-(k_{1}+k_{2})}{k_{3},k_{4},\dots,k_{m}}$. (20)
@@
$\sum_{\begin{subarray}{c}0\leq k_{1},k_{2},\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}$
$\ =\ \sum_{\begin{subarray}{c}0\leq k_{1},k_{2}<N^{\epsilon}\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}N^{k_{1}+k_{2}}\sum_{\begin{subarray}{c}0\leq k_{3},k_{4},\dots,k_{m}\leq N-(k_{1}+k_{2})\\ k_{3}+k_{4}+\cdots+k_{m}=N-(k_{1}+k_{2})\end{subarray}}\binom{N-(k_{1}+k_{2})}{k_{3},k_{4},\dots,k_{m}}$
@@
$\ \leq\ (m-2)^{N}(N^{\epsilon})^{2}N^{N^{\epsilon}}\ \leq\ (m-2)^{N}N^{N^{\epsilon}+2}.$ (21)
@@
$\frac{1}{m^{N}}\sum_{\begin{subarray}{c}0\leq k_{1},k_{2}\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}$. (22)
According to Proposition 2, we have that (22) is bounded above by $((m-2)/m)^{N}N^{N^{\epsilon}+2}$. Hence, the logarithm of the probability is bounded above by
$\log\left(\left(\frac{m-2}{m}\right)^{N}N^{N^{\epsilon}+2}\right)\ =\ N\log\left(\frac{m-2}{m}\right)+(N^{\epsilon}+2)\log(N).$ (23)
It is clear that the above goes to -$\infty$ as $N\rightarrow\infty$. Thus the probability goes to 0 as $N\rightarrow\infty$. So it suffices to consider the case where $k_{1}+k_{2}\geq N^{\epsilon}$. Now, fix $k(N)\geq N^{\epsilon}$ and let $0\leq k_{1},k_{2}\leq k(N)$ with $k_{1}+k_{2}=k(N)$ and $0\leq k_{3},k_{4},\dots,k_{m}\leq N-k(N)$ with $k_{3}+k_{4}+\dots+k_{m}=N-k(N)$. By [1], we know that the frequency $k_{1}$ follows a binomial distribution with mean $k(N)/2$ and standard deviation $\sqrt{k(N)}/2$. Pick some $\delta\in(0,\epsilon/10)$. We see that it suffices to consider cases where $|k_{1}-k(N)/2|<(\lceil N^{\delta}\rceil\sqrt{k(N)})/2$, because the probability that $k_{1}$ is outside this range is asymptotically small by Chebyshev’s inequality:
$\mathbb{P}\left(\left|k_{1}-\frac{k(N)}{2}\right|\ \geq\ \frac{\lceil N^{\delta}\rceil\sqrt{k(N)}}{2}\right)\ \leq\ \frac{1}{\lceil N^{\delta}\rceil^{2}}.$ (24)
2.4 Near uniform probability within small intervals
@@
$k_{1,\ell}\ :=\ f_{\ell}\left(\frac{k(N)}{2}\right)+\ell\lceil N^{\delta}\rceil,$
$k_{1,\ell,i}\ :=\ f_{\ell}\left(\frac{k(N)}{2}\right)+\ell\lceil N^{\delta}\rceil+i,\quad 0\leq i\leq N^{\delta}-1,$ (25)
@@
We follow Section (5.52) of [1] to obtain a bound for the difference.
Proposition 3.
@@
$\left|\binom{k(N)}{k_{1,\ell}}-\binom{k(N)}{k_{1,\ell+1}}\right|\ \leq\ O\left(\binom{k(N)}{k_{1,\ell}}\cdot N^{-\frac{3\epsilon}{10}}\right).$ (27)
Since the proof follows a similar argument to Section (5.52) of [1], we leave the proof details to A.
2.5 Equidistribution within small intervals
In this subsection, we want to show that for fixed $k(N),k_{3},\dots,k_{m}$, the logarithm of the stick length $\log_{B}(A^{(N)}_{k_{1,\ell,i},k_{2},\dots,k_{m}})$ for $0\leq i\leq\lceil N^{\delta}\rceil-1$ converges to being equidistributed mod 1, in the sense that the logarithm of the stick length is thought of as a random variable that takes value in $\log_{B}(A^{(N)}_{k_{1,\ell,i},k_{2},\dots,k_{m}})$ for $0\leq i\leq\lceil N^{\delta}\rceil-1$ with equal probabilities. We first state the following theorem which provides an easy criterion for checking equidistribution mod 1.
Corollary 1.
[15] For any $a\in\mathbb{R}$, the arithmetic progression $\{a+nd\}_{n=1}^{\infty}$ is equidistributed mod 1 if and only if $d\not\in\mathbb{Q}$.
Now, note that for fixed $k(N)$ and $\ell$,
$\log_{B}(A^{(N)}_{k_{1,\ell,i},k_{2},\dots,k_{m}})$
$\ =\ k_{1,\ell,i}\log_{B}\left(\frac{p_{1}}{p_{2}}\right)+(k_{1}+k_{2})\log_{B}\left(\frac{p_{2}}{p_{3}}\right)+\cdots+\left(\sum_{j=1}^{m-1}k_{j}\right)\log_{B}\left(\frac{p_{m-1}}{p_{m}}\right)+N\log(p_{m})$
$\ =\ k_{1,\ell,i}\log_{B}\left(\frac{p_{1}}{p_{2}}\right)+k(N)\log_{B}\left(\frac{p_{2}}{p_{3}}\right)+\cdots+\left(\sum_{j=1}^{m-1}k_{j}\right)\log_{B}\left(\frac{p_{m-1}}{p_{m}}\right)+N\log(p_{m})$ (28)
@@ 2.5 Equidistribution within small intervals
Definition 10.
[15] Suppose that $x$ is a real number. The irrationality exponent $\mu$ of $x$ is the supremum of the set of $\mu$ such that $0<|x-p/q|<1/q^{\mu}$ is satisfied by an infinite number of coprime integer pairs $(p,q)$ with $q>0$. If such a set does not exist, then we say $x$ has irrationality exponent $\infty$.
Since $\mathbb{Q}$ is dense in $\mathbb{R}$, then every real number can be approximated by rational numbers. However, what differentiates them is how well they can be approximated. Irrationality exponent measures exactly how well a real number can be approximated by rational numbers. The bigger the irrationality exponent, the finer the rational approximation can be. The following are some well-known facts about irrationality exponent, which comes handy later in this section.
Proposition 4.
[15] The irrationality exponent of any rational number is 1. As a consequence of Dirichlet’s approximation theorem, which states that for any irrational number $x$,
$\left|x-\frac{p}{q}\right|\ <\ \frac{1}{q^{2}}$ (29)
for infinitely $(p,q)$ where $p$ and $q$ are coprime, the irrationality exponent of any irrational number is at least 2. Further more, any algebraic irrational (irrational numbers that are zeros of polynomials over $\mathbb{Q}$) has irrationality exponent exactly 2.
We now state a theorem that quantifies the discrepancy of the logarithm of the stick length in terms of irrationality exponent.
Theorem 6.
[15, Theorem 3.2] Let $\kappa$ be the irrationality exponent of $\log_{B}(p_{1}/p_{2})$ and $J_{\ell}(a,b)\subset\{0,1,\dots,N^{\delta}-1\}$ be the set of indices $i$ such that $\log_{B}(A^{(N)}_{k_{1,\ell,i},k_{2},\dots,k_{m}})$ mod 1 is in $(a,b)$. If $\kappa<\infty$, then
$|J_{\ell}(a,b)|\ =\ (b-a)N^{\delta}+O\left(N^{\delta\left(1-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\right),$ (30)
for every $\epsilon^{\prime}>0$ and there is a power saving with the error term. The error term is optimal in the sense that
$\left|\frac{|J_{\ell}(a,b)|}{N^{\delta}}-(b-a)\right|\ =\ \Omega\left(N^{\delta(-\frac{1}{\kappa-1})-\epsilon^{\prime}}\right)$ (31)
for every $\epsilon^{\prime}>0$. If $\kappa=\infty$, then
$|J_{\ell}(a,b)|\ =\ (b-a)N^{\delta}+o(N^{\delta}).$ (32)
Note that [15] has a different convention for irrationality exponent and if the irrationality exponent of $x$ is $\mu$ in our definition, then it is $\mu-1$ in their definition. This is the reason for the change made to the presentation of Theorem [15], where there is a $1/(\kappa-1)$ instead of a $1/\kappa$ in the exponent.
2.6 Evaluation of sums within and across intervals